OpenStudy (anonymous):

Find G'(x) if G(x) (integral with a to x) 2lnydy???

OpenStudy (anonymous):

\[G(x) \int\limits_{a}^{x}2lnydy \]

myininaya (myininaya):

\[G(x)=\int\limits\limits_{a}^{x}f(y) dy = F(y)|_{a}^{x} =F(x)-F(a) \] F(a) is a constant so ... \[(G(x))'=(F(x)-F(a))'=(F(x))' \] But F'=f so we have \[G'(x)=F'(x)=f(x)\]

OpenStudy (anonymous):

so what would I put in for the x and a?

myininaya (myininaya):

G'(x)=f(x) where f is the machine inside your integral just replace the y with x.

myininaya (myininaya):

Why do we want to put something there to replace them? We are trying to find the derivative of the above expression.

OpenStudy (anonymous):

so the answer would be 2lnxdx?

myininaya (myininaya):

not the dx

OpenStudy (anonymous):

So just 2lnx? and thats it?

myininaya (myininaya):

yep

OpenStudy (anonymous):

would it look like \[\int\limits_{a}^{x}(2lnx)dx?\]

myininaya (myininaya):

?

myininaya (myininaya):

No we were finding the derivative of G

OpenStudy (anonymous):

So just 2lnx? when I type it into my homework it makes it look like 2ln(x), is that okay?

myininaya (myininaya):

Yes. Did you not understand the first post? If you don't please tell me what you don't understand so I can better help you.

OpenStudy (anonymous):

I think it just confused me about the parenthesis, but if its okay as 2ln(x) than I think I understand! Thanks for your help it was right!! :)